Operation Summary for February 21, 2011 - February 23, 2011 |
|
Monday Feb. 21 The day shift began with Operations monitoring the Tevatron (TEV) store 8518 (55.03E30), the Recycler stash (339.8E10), and the Antiproton Source (Pbar) stack (9.6E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations terminated store 8518 at 8:20 AM and established store 8519 at 10:52 AM with an average initial luminosity (AIL) of 302E30. At noon, operators turned off Pbar for repairs to its target station LCW system. During the evening shift, operators monitored the accelerator complex. Tuesday Feb. 22 The midnight shift began with Operations monitoring store 8519 (82.16E30) and the stash (71.5E10), with no Pbar stack due to repairs, and with NuMI and MiniBooNE taking beam. During the midnight shift, at 1:45 AM, a fault in the NuMI humidity control system halted beam for 2.25 hours. During the day shift, Operations terminated store 8519. TEV personnel conducted studies throughout the day shift. The Pbar LCW work concluded by 10:30 AM. Pbar resumed stacking. During the evening shift, TEV studies continued. Operations established store 8526 at 9:43 PM with an AIL of 272E30. Wednesday Feb. 23 The midnight shift began with Operations monitoring store 8526 (168.2E30) and the stash (86.3E10), with Pbar stacking (15.5E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, operators monitored the accelerator complex. The Plans for Wednesday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 02/23/2011 email Fermilab |
| Security, Privacy, Legal | |